If a third degree polynomial has a lone x-intercept at

x = a , discuss what this implies about the

linear and quadratic factors of that polynomial.


1
Expert's answer
2021-10-20T05:46:52-0400

Since a third degree polynomial "f" has a lone "x"-intercept at "x=a", it has a unique root "x=a." If the multiplicity of this root is 3, then "f(x)=c(x-a)^3," and in this case the polynomial has three (the same) linear factors. If the root "x=a" is of multiplicity 2, then the third factor is linear, and polinomial has the root "x=b\ne a," which is impossible according to uniqueness of a root. If the multiplicity of this root is 1, then "f(x)=(x-a)(bx^2+cx+d)," and polynomial "g(x)=bx^2+cx+d" has no roots. It follows that the last polynomial is irreducible, and hence it is the quadratic factor of polynomial "f".